16. What volume will 12 g of oxygen gas (O2) occupy at 25 °C and a pressure of 52.7 kPa?



1
Expert's answer
2021-04-25T07:39:24-0400

Ideal gas law equation "PV=nRT"


"P-pressure =52.7kPa"

"V-volume=?"

"n-moles=\frac {12}{32}=0.375moles"

"R=Constant = 8.314Jmol^{-1}K^{-1}"

"T=Temperature = 25 +273=298K"


Substituting the values

"52700Pa\u00d7V=0.375mol\u00d78.314Jmol^{-1}K^{-1}\u00d7298K"


"V=17.6L"